OpenStudy (anonymous):

Write the equation in slope-intercept form. 5x - y =7 and 2x + 5y= 15 Then tell me how you do it?

OpenStudy (anonymous):

Do you know what slope intercept form looks like?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

Perfect, so, all you're going to do is move the symbols/numbers around so it looks the same as that.

OpenStudy (anonymous):

5x-y=7 We will leave y on the left because that's where we want it. Subtract 5x from both sides: -y=7-5x Divide both sides by -1 to get rid of the - on the y: y=-7+5x Switch the places of the -7 and 5x so it looks like the form you gave: y=5x-7 m=5 and b=-7

OpenStudy (anonymous):

Then what would the second one be?

OpenStudy (anonymous):

Give it a shot and I'll help you work through it

OpenStudy (anonymous):

I can't figurte this one out, I think it's like y=5 then I get lost

OpenStudy (anonymous):

2x + 5y= 15 Do it one step at a time. Try to get y by itself. You can't divide both sides by 5 to get rid of the 5 next to the y until you get rid of that 2x. What can you do to get rid of the 2x on the left?

OpenStudy (anonymous):

Divide by twp on both sides?

OpenStudy (anonymous):

Well, you want to get rid of the whole 2x not just 2. So, look at that as a chunk by itself. 2x is being added to 5y. What's the opposite of addition?

OpenStudy (anonymous):

subtraction

OpenStudy (anonymous):

So, subtract 2x from both sides. What do you get?

OpenStudy (phi):

Here is a short video that shows how to do this http://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/v/converting-to-slope-intercept-form